OpenStudy (anonymous):

Match the following equation to the correct situation. 60%(p + 3.2) = 60%p + 1.92

OpenStudy (anonymous):

60% is the same thing as 0.6 \(0.6(p + 3.2) = 0.6p + 1.92\) Distribute 0.6 into the parenthesis: \(0.6p + 1.92 = 0.6p + 1.92\) Notice anything?

OpenStudy (anonymous):

Look at the left of the equal sign and to the right of the equal sign.

OpenStudy (anonymous):

@iGreen. Sorry I was eating.

OpenStudy (anonymous):

No problem, what do you see?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Um, yes, I distributed, but what do you see about the two different sides of the equal sign?

OpenStudy (anonymous):

You added 1.92

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

All I did was multiply: 0.6 * p 0.6 * 3.2 To get rid of the parenthesis.

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

What are your choices?

OpenStudy (anonymous):

0.6p + 1.92 = 0.6p + 1.92 Both sides of the equal sign are the same.

OpenStudy (anonymous):

UGHHH, my session xpired

OpenStudy (anonymous):

? Now what?

OpenStudy (anonymous):

I can't go back to that question.

OpenStudy (anonymous):

Oh..so I guess you should close this..